OpenStudy (anonymous):

A rectangle is 2 times as long as it is wide. If the area is 8 square feet, find its perimeter.

OpenStudy (amistre64):

8 = 2ww

OpenStudy (anonymous):

2w = l w * l = 8 w * 2w = 8 w*w = 4 w = 2 l = 4 p = 2(2) + 2(4) p = 12